OpenStudy (sambam):

The line with m = 2/3 and passing through (1, 1)

OpenStudy (anonymous):

The slope-intercept form for the equation of a line is\[y-y _{1}=m(x-x _{1})\] when the point is \[(x _{1}, y _{1})\] Now plug in your points and your slope to get \[y-1=\frac{ 2 }{ 3 }(x-1)\]

OpenStudy (anonymous):

Then distribute, simplify, and solve for y \[y-1=\frac{ 2 }{ 3 }x-\frac{ 2 }{ 3 }\] \[y=\frac{ 2 }{ 3 }x-\frac{ 2 }{ 3 }+1\] \[y=\frac{ 2 }{ 3 }x+\frac{ 1 }{ 3 }\]